OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form ?

OpenStudy (anonymous):

slope of line perp. to this one will be -1/4( cause product of perp. lines is -1) use the general eq. of line passing through a given point y - a = m(x-b) ((a,b) is the point) y-(-3) = -1/4(x-2) y+3 = -1/4x + 1/2 y + 1/4x = -5/2 4y + x = -10 on multiplying the equation be 4

OpenStudy (anonymous):

THANK YOU! Now i get it